# Leveraging Auxiliary Task Relevance for Enhanced Industrial Fault Diagnosis through Curriculum Meta-learning

### v1

@@ Abstract
The accurate diagnosis of machine breakdowns is crucial for maintaining operational safety in smart manufacturing. Despite the promise shown by deep learning in automating fault identification, the scarcity of labeled training data, particularly for equipment failure instances, poses a significant challenge. This limitation hampers the development of robust classification models. Existing methods like model-agnostic meta-learning (MAML) do not adequately address variable working conditions, affecting knowledge transfer. To address these challenges, a Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed in this paper, inspired by human cognitive learning processes. RT-ACM improves training by considering the relevance of auxiliary working conditions, adhering to the principle of “paying more attention to more relevant knowledge”, and focusing on “easier first, harder later” curriculum sampling. This approach aids the meta-learner in achieving a superior convergence state. Extensive experiments on two real-world datasets demonstrate the superiority of RT-ACM framework.
@@ I Introduction
Fault diagnosis is crucial for ensuring the safe and efficient operation of machines [1]. Recently, methods based on deep learning (DL) have demonstrated considerable promise in accurately diagnosing machine faults [2]. An advanced ensemble convolutional neural network (CNN) has been introduced for this purpose, utilizing diverse local minima to enhance model adaptability [3]. Furthermore, a capsule network that builds upon CNNs has been proposed for fault classification in bearings [4]. This network combines an inception block and a regression branch, thereby increasing the model’s versatility. However, a notable challenge faced by DL methods is the requirement for extensive training data. Accumulating such data in real-world scenarios is often not feasible due to associated costs and logistical challenges [5]. When machines exhibit faults, they are typically partially (e.g., showed down) or fully (e.g., turned off) deactivated for safety reasons, resulting in a scarcity and variety of fault data [6]. Although intentional fault inductions on lab-based machines might provide data, such interventions in operational settings are strictly prohibited. This situation intensifies the challenge of developing reliable DL classifiers for fault diagnosis in real-world contexts [7].
@@
Three key strategies have emerged to tackle the challenge of data scarcity in creating reliable fault diagnosis models: Data Augmentation (DA) [8], Transfer Learning (TL) [9], and Few-Shot Learning (FSL) approaches [10]. At its essence, DA aims to expand and diversify datasets by producing additional data from the limited original samples available. For example, data enhancement techniques that incorporate Gaussian noise, masking noise, and signal translation have been explored in [8]. Moreover, generative models, including variational autoencoders and generative adversarial networks, have been employed to generate fault data for training DL-based diagnostic models in [11]. However, a notable limitation of DA is the potential degradation in data quality and precision, especially when original datasets are sparse. This might reduce the effectiveness of models trained on such data.
@@
To circumvent the limitations of DA, the use of prior knowledge from related tasks has been explored, leading to the concept of Transfer Learning (TL). TL methods strive to utilize knowledge derived from one task to benefit a different, albeit related, task. For instance, a comprehensive one-dimensional CNN architecture, supplemented with fine-tuning mechanisms specifically for fault diagnosis, has been presented[1]. Similarly, a CNN-based model for diagnosing faults in bearings has been proposed, with its basis being knowledge garnered from laboratory bearings data[12]. However, a potential drawback of TL techniques is their inherent emphasis on adapting models to target tasks, which may hinder their broad applicability, especially in environments with limited data [13]. Due to this constraint, there exists a growing need to enhance the overall generalization capabilities of TL methods for more effective fault diagnosis in situations characterized by data sparsity.
@@
Few-shot learning (FSL) stands out for its capability to generalize across tasks, not merely adapting to one. It achieves this by leveraging insights from various auxiliary tasks, a method commonly known as meta-learning [14]. This stands in contrast to transfer learning, where the primary focus is adaptation to a distinct task. FSL delves deeper by leveraging insights from associated tasks to benefit the primary FSL task [15]. In practical settings, machines typically operate under varied conditions. Gathering substantial data for each specific condition is both expensive and, in some cases, impractical. Yet, when data from varied operational conditions are utilized, a multitude of auxiliary tasks can be formulated. This inherent flexibility renders meta-learning especially fitting and potent for few-shot fault diagnosis scenarios. Such FSL models specifically designed for fault diagnosis have been proposed in [16]. These models frequently incorporate metric-based meta-learning methods, with Siamese neural networks and matching networks being prominent examples.
@@
Despite the potential of metric-based meta-learning methods for feature extraction, their efficacy is hindered by data sparsity and variations in auxiliary task distributions. Model-Agnostic Meta-Learning (MAML) [17], a optimization-based meta-learning method, diverges from traditional metric-based approaches by emphasizing the development of models with robust generalization capabilities. MAML aims to find an initial parameters set, then the model could achieve optimal performance on a series of new tasks after one or more gradient updates, which are calculated using a small amount of data from the series of tasks. Through minimal fine-tuning with target task data, MAML demonstrates superior classification performance in few-shot bearing fault diagnosis. Zhang et al. [18] proposed a MAML-based few-shot learning framework for bearing fault diagnosis with limited data. They achieved this by developing a fault diagnosis model using an improved meta-relation network. Zhou et al. [19] introduced an augmented method that integrates prior knowledge into MAML to fully utilize the prior knowledge present in both training and testing sets. Chen et al. [10] proposed a MAML based model featuring a four-layer CNN capable of feature extraction and rapid adaptation, enabling it to quickly adapt to fault diagnosis tasks under various working conditions. Although these studies have achieved some performance improvements and emphasized that MAML learns to learn, they have overlooked the difficulty level of each task and have not considered the human learning experience of easy-to-hard curriculum learning. Chen et al. [20] propose a novel curriculum hardness aware meta-Learning framework, which incorporates hard sample mining and curriculum learning into a meta-learning paradigm. Hu et al. [6] utilized a specialized task-sequencing strategy for few-shot fault diagnosis, introducing a method where auxiliary tasks are arranged from simple to complex to improve knowledge adaptability during meta-learning. Nevertheless, these existing studies have not considered the relevance between different tasks and the target task. Irrelevant auxiliary tasks may negatively impact the target task.
@@
To simultaneously consider the human easy-to-hard curriculum learning processing and the relevance between different tasks and the target task, we propose a novel Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, which delicately considers the relevance of patterns among different tasks within the meta-learning paradigm to alleviate data sparsity and task diversity issues. To begin with, RT-ACM extends MAML to facilitate knowledge transfer to the target task under conditions of limited data. With fault diagnosis in different working conditions treated as single tasks, the initial model weights are meta-learned for faster adaptation to limited training data in target tasks. Furthermore, we design two components in the meta-training process: related task-aware meta-learning and a curriculum strategy to explicitly address the task diversity issue.
@@
Firstly, we exploit the heuristic idea of “pay more attention to more relevant knowledge” to enhance the meta-learner by considering the relevance of auxiliary tasks as related task-aware meta-learning [21]. This approach compels the model to learn more from a combination of tasks that are more similar and are expected to be highly diverse, thereby contributing more to generalization. Specifically, we divide each meta-training round into two stages. In each stage, we first update the meta-learner in a MAML-like manner and then conditionally re-sample a new batch of tasks, focusing on the most complex working conditions based on validation score. Secondly, we propose drawing upon a curriculum strategy to improve the convergence rate and generalization capacity of the meta-learner given the high diversity among different tasks. The underlying concept is to present training tasks to the meta-learner in an “easier first, harder later” paradigm as determined by a teacher recommender [20, 22]. This aims to guide the meta-learner towards a better state. Specifically, we pre-train a teacher recommender for each auxiliary task, use the best validation score to gauge task difficulty, and utilize this knowledge to design a curriculum for the task sampling pool at each meta-training step.
Overall, our main contributions lie in three folds:
@@
    To the best of our knowledge, it is one of the early studies that examines to what extent knowledge can be transferred from auxiliary tasks to the target task by differentiating the relevance of working conditions.
@@
    A novel Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed, which leverages both transferred knowledge and the diversity of tasks within the target task to mitigate the issue of data sparsity.
@@
    Extensive experiments have been conducted on two real-world datasets to validate the superiority of RT-ACM against state-of-the-art approaches.
The rest of this article is organized as follows. The proposed method is detailed in Section II. In Section III, the effectiveness of RT-ACM is evaluated with various case studies. Finally, Section IV concludes this article and presents future work.
@@ II Proposed RT-ACM
Assume that $A$ auxiliary tasks are available and denoted by $\left\{T^{\mu}\right\}_{\mu=1}^{A}$. In each auxiliary task, $T^{\mu}=\left\{(x^{\mu}_{i},y^{\mu}_{i})\right\}^{M}_{i=1}$, where $x^{\mu}_{i}\in\mathcal{R}^{1\times D}$ is $i$th sample of the input data. In the field of bearing fault diagnosis, 1-D vibration signal is commonly used as the input data. The vibration signal is segmented using a sliding window to generate samples, and $D$ is the length of the sliding window. $y_{i}^{\mu}$ is the health status of $x_{i}^{\mu}$. $M$ denotes the number of samples contained in $T^{\mu}$.
A target task is denote by $T^{tar}=\left\{(x^{tar}_{i},y^{tar}_{i})\right\}^{M^{tar}}_{i=1}$, where only a few samples are available for training. Thus, $M^{tar}$, the number of samples contained in $T^{tar}$, is assigned a very small value. $x^{tar}_{i}\in\mathcal{R}^{1\times D}$ is the input data of the $i$th sample and $y^{tar}_{i}$ is the corresponding label.
The overview of RT-ACM is outlined in Fig. 1, mainly composed of two channels (i.e., meta-level training on auxiliary tasks and target-level fine-tune). In particular, the meta-level training exploits meta-learning to capture the common vibration pattern by holding the principle of “paying more attention to more relevant knowledge” and learning knowledge curriculum judge by teacher recommender. The goal of target-level fine-tune is to learn the accurate vibration pattern in the target task and performs the predictions.
@@ II-A Meta-level Training on Auxiliary Tasks
To distil knowledge from auxiliary tasks and employ vibration patterns, we extend the MAML with long short-term memory (LSTM) [23] as the framework for the meta-learning update. In particular, we devise a related task aware meta-learning and curriculum strategy that can transfer knowledge based on the relevance of patterns among tasks.
@@ II-A1 Basic Theory of MAML
Unlike traditional deep learning, as a category of meta-learning, the MAML algorithm is designed to train the initialization parameters of a model within the outer loop. By learning a set of initialization parameters that are sufficiently general, the model can achieve robust performance efficiently with only a few gradient update steps when adapting to new tasks with limited data. From a learning perspective, MAML seeks to identify general internal representations that facilitate transferability to various tasks. Consequently, when confronted with new tasks, only minimal parameter fine-tuning is necessary.
Following [20], the fault diagnosis in different working conditions can be viewed as a single task in a meta-learning paradigm. Thus, the vibration signal of auxiliary tasks $\left\{T^{\mu}\right\}_{\mu=1}^{A}$ are denoted as $\mathbb{D}_{meta}^{(aux)}$ and the vibration signal of target tasks $T^{tar}$ are divided as training sets $\mathbb{D}_{train}^{(tar)}$ and test sets $\mathbb{D}_{test}^{(tar)}$. We treat each working conditions $y_{m}$ as a meta-learning task, where each task has support set $\mathcal{D}^{spt}_{y_{m}}$ for training and a query set $\mathcal{D}^{qry}_{y_{m}}$ for testing. Finally, our goal is to leverage the auxiliary tasks to learn a meta-learner $F_{w}$, which contains knowledge about different working conditions.
Specifically, each iteration of MAML includes local update and global update on the sampled task batch, where the first phase updates $\theta$ locally on the $\mathcal{D}^{spt}$ of each task, and the second phase globally updates $\theta$ by gradient descent to minimize the sum of loss on the $\mathcal{D}^{qry}$ of all tasks.
  • Local update: we first sample a batch of tasks from $\mathbb{D}^{(aux)}_{meta}$. Then we randomly sample a group of vibration signal from $\mathcal{D}^{spt}_{y_{m}}$ and $\mathcal{D}^{qry}_{y_{m}}$ for each sampled tasks. Thus, we calculate the training loss on $\mathcal{D}^{spt}_{y_{m}}$ and locally update $\theta$ by one step:
@@
  • Global update: we calculate the testing loss on each $\mathcal{D}^{qry}_{y_{m}}$ with the corresponding $\theta_{y_{m}}^{\prime}$. The MAML global updating aims to update the initialization $\theta$ by one gradient step on the sum of all the testing losses:
@@
The optimal parameters of the meta-learning model can be obtained through alternating optimization of the inner and outer loops. Therefore, the goal of the meta-optimization is to minimize the loss function of the task:

### v2

# Leveraging Auxiliary Task Relevance for Enhanced Bearing Fault Diagnosis through Curriculum Meta-learning

@@ Abstract
The accurate diagnosis of machine breakdowns is crucial for maintaining operational safety in smart manufacturing. Despite the promise shown by deep learning in automating fault identification, the scarcity of labeled training data, particularly for equipment failure instances, poses a significant challenge. This limitation hampers the development of robust classification models. Existing methods like model-agnostic meta-learning (MAML) do not adequately address variable working conditions, affecting knowledge transfer. To address these challenges, a Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed in this paper, inspired by human cognitive learning processes. RT-ACM improves training by considering the relevance of auxiliary sensor working conditions, adhering to the principle of “paying more attention to more relevant knowledge”, and focusing on “easier first, harder later” curriculum sampling. This approach aids the meta-learner in achieving a superior convergence state. Extensive experiments on two real-world datasets demonstrate the superiority of RT-ACM framework.
@@ I Introduction
Fault diagnosis is crucial for ensuring the safe and efficient operation of machines [1, 2]. Recently, methods based on deep learning (DL) have demonstrated considerable promise in accurately diagnosing machine faults through sensor signals [3, 4]. An advanced ensemble convolutional neural network (CNN) has been introduced for this purpose, utilizing diverse local minima to enhance model adaptability [5]. Furthermore, a capsule network that builds upon CNNs has been proposed for fault classification in bearings [6]. This network combines an inception block and a regression branch, thereby increasing the model’s versatility. However, a notable challenge faced by DL methods is the requirement for extensive training data. Accumulating such data in real-world scenarios is often not feasible due to associated costs and logistical challenges [7]. When machines exhibit faults, they are typically partially (e.g., showed down) or fully (e.g., turned off) deactivated for safety reasons, resulting in a scarcity and variety of fault data [8]. Although intentional fault inductions on lab-based machines might provide data, such interventions in operational settings are strictly prohibited. This situation intensifies the challenge of developing reliable DL classifiers for fault diagnosis in real-world contexts [9].
@@
Three key strategies have emerged to tackle the challenge of data scarcity in creating reliable fault diagnosis models: Data Augmentation (DA) [10], Transfer Learning (TL) [11], and Few-Shot Learning (FSL) approaches [12]. DA aims to expand and diversify datasets by producing additional data from the limited original samples available. For example, data enhancement techniques that incorporate Gaussian noise, masking noise, and signal translation have been explored in [10, 13]. Moreover, generative models, including variational autoencoders and generative adversarial networks, have been employed to generate sensor fault data for training DL-based diagnostic models in [14]. However, a notable limitation of DA is the potential degradation in data quality and precision, especially when original datasets are sparse. This might reduce the effectiveness of models trained on such data.
@@
To circumvent the limitations of DA, the use of prior knowledge from related tasks has been explored, leading to the concept of Transfer Learning (TL). TL methods aim to transfer knowledge gained from one task to improve performance on another closely related task. For instance, a comprehensive one-dimensional CNN architecture, supplemented with fine-tuning mechanisms specifically for fault diagnosis, has been presented[1, 15]. Similarly, a CNN-based model for diagnosing faults in bearings has been proposed, with its basis being knowledge garnered from laboratory bearings data[16]. However, a potential drawback of TL techniques is their inherent emphasis on adapting models to target tasks, which may hinder their broad applicability, especially in environments with limited data [17]. Due to this constraint, there exists a growing need to enhance the overall generalization capabilities of TL methods for more effective fault diagnosis in situations characterized by data sparsity.
@@
Few-shot learning (FSL) stands out for its capability to generalize across tasks, not merely adapting to one. It achieves this by leveraging insights from various auxiliary tasks, a method commonly known as meta-learning [18]. This stands in contrast to transfer learning, where the primary focus is adaptation to a distinct task. FSL delves deeper by leveraging insights from associated tasks to benefit the primary FSL task [19]. In practical settings, machines typically operate under varied conditions, gathering substantial sensor data for each specific condition is both expensive and, in some cases, impractical [20]. Yet, when data from varied operational conditions are utilized, a multitude of auxiliary tasks can be formulated. This inherent flexibility renders meta-learning especially fitting and potent for few-shot fault diagnosis scenarios. Such FSL models specifically designed for fault diagnosis have been proposed in [21, 22]. These models frequently incorporate metric-based meta-learning methods.
@@
Metric-based meta-learning methods have demonstrated potential in feature extraction; however, their effectiveness is often limited by data sparsity and variations in auxiliary task distributions. Model-Agnostic Meta-Learning (MAML) [23], an optimization-based meta-learning framework, offers a different approach by focusing on developing models with strong generalization capabilities. MAML seeks to identify a set of initial parameters that can be fine-tuned with minimal data to achieve optimal performance across a series of new tasks. As shown in Fig. 1, MAML optimizes the model to establish a representation, $\theta$, that facilitates rapid adaptation to new tasks. By leveraging a small amount of target task data for fine-tuning, MAML achieves notable classification accuracy in few-shot bearing fault diagnosis. For instance, Zhang et al. [24] proposed a MAML-based framework employing an enhanced meta-relation network for bearing fault diagnosis with limited data. Zhou et al. [25] further advanced this concept by integrating prior knowledge into MAML, enabling more effective utilization of prior knowledge in both training and testing phases. Chen et al. [12] developed a MAML-based model featuring a four-layer CNN designed for feature extraction and rapid adaptation, allowing efficient fault diagnosis under diverse working conditions. Additionally, Cui et al. [26] employed a federated meta-learning approach combining representation encoding with MAML to enhance performance in distributed fault diagnosis tasks. Although these studies have achieved some performance improvements and emphasized that MAML learns to learn, they have overlooked the difficulty level of each task and have not considered the human learning experience of easy-to-hard curriculum learning. A specialized task-sequencing strategy, task-sequencing meta-learning (TSML) [8], has been introduced for few-shot fault diagnosis. This approach arranges auxiliary tasks from simple to complex, aiming to enhance knowledge adaptability during meta-learning. However, it relies on k-means clustering to evaluate task difficulty, which overlooks the intricate relationships within temporal data, limiting its effectiveness to simpler datasets. Furthermore, these existing studies have not considered the relevance between different tasks and the target task, as unrelated auxiliary tasks can adversely affect the target task.
Figure 1: Learning process of MAML, which optimizes for a representation $\theta$ that can quickly adapt to new tasks.
@@
To simultaneously consider data sparsity and task diversity challenges in fault diagnosis, we propose a novel Related Task Aware Curriculum Meta-learning (RT-ACM) framework that integrates human-inspired easy-to-hard curriculum learning with task relevance considerations. RT-ACM leverages the meta-learning paradigm to capture the relationships among patterns from diverse tasks, enabling more effective knowledge transfer under limited data conditions. Building on the foundation of MAML, the framework is designed to enhance adaptation to target tasks with minimal training data by treating fault diagnosis tasks under varying working conditions as individual tasks. The meta-training process incorporates two key components: a related task-aware meta-learning mechanism that emphasizes task relevance, and a curriculum strategy that systematically addresses task diversity, facilitating improved learning efficiency and generalization.
@@
To enhance the meta-learner, we exploit the heuristic idea of ”paying more attention to more relevant knowledge,” adopting a related task-aware meta-learning approach. This strategy encourages the model to prioritize learning from tasks that are highly relevant, thereby improving generalization. Each meta-training round is structured into two stages: first, the meta-learner is updated using a MAML-like process; second, a new batch of tasks is conditionally re-sampled, focusing on the most challenging working conditions as determined by validation scores. Specifically, we propose a curriculum strategy to refine the task sampling process during MAML training. Guided by the principle of an ”easier-first, harder-later” paradigm, this approach directs the meta-learner toward an optimized state. A teacher recommender is pre-trained for each auxiliary task, leveraging the best validation score to assess task difficulty and dynamically designing the task sampling curriculum at each meta-training step.
Overall, our main contributions can be summarized in three key aspects:
@@
    To the best of our knowledge, this is one of the early studies that examines the extent to which knowledge can be transferred from auxiliary tasks to the target task by differentiating the relevance of working conditions. A novel related task-aware curriculum meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed, which leverages both transferred knowledge and the diversity of tasks within the target task to mitigate the issue of data limitation.
@@
    A meta-training process is formulated, which incorporates related task-aware meta-learning and a curriculum strategy to leverage auxiliary tasks while balancing the high diversity among different tasks. The related task-aware meta-learning follows the principle of “paying more attention to more relevant tasks”, while the curriculum strategy follows the principle of “easy first, hard later.” Moreover, we design a teacher recommender to evaluate task difficulty in temporal data.
@@
    Extensive experiments have been conducted on two real-world datasets to validate the superiority of RT-ACM against state-of-the-art approaches. Furthermore, we have made our source code publicly available to contribute further to advancements in this field.
The remainder of this article is structured as follows: Section II provides a detailed description of the proposed method. Section III presents a series of case studies to evaluate the effectiveness of RT-ACM. Finally, Section IV offers conclusions and outlines directions for future work.
@@ II Proposed RT-ACM
Let $A$ auxiliary tasks be represented as $\left\{T^{\mu}\right\}_{\mu=1}^{A}$, where each auxiliary task $T^{\mu}$ comprises a dataset $\left\{(x^{\mu}_{i},y^{\mu}_{i})\right\}_{i=1}^{M}$. Here, $x^{\mu}_{i}\in\mathcal{R}^{1\times D}$ corresponds to the $i$th input sample derived from 1-D vibration sensor signals, a common data source in this domain. These signals are segmented using a sliding window of length $D$ to produce samples, and $y^{\mu}_{i}$ denotes the sensor health status label associated with $x^{\mu}_{i}$. The total number of samples in $T^{\mu}$ is given by $M$. Similarly, a target task is defined as $T^{tar}=\left\{(x^{tar}_{i},y^{tar}_{i})\right\}_{i=1}^{M^{tar}}$, where the number of available training samples, $M^{tar}$, is limited, reflecting the scarcity of labeled data. Each $x^{tar}_{i}\in\mathcal{R}^{1\times D}$ represents the input data for the $i$th sample, while $y^{tar}_{i}$ is its corresponding label.
The structure of RT-ACM is depicted in Fig. 2, consisting primarily of two components: meta-level training on auxiliary tasks and fine-tuning on the target task. Specifically, the meta-level training employs meta-learning to extract shared sensor vibration patterns, adhering to the principles of “paying more attention on more relevant knowledge” and “easy first, hard later,” as determined by the teacher recommender. Meanwhile, the target-level fine-tuning aims to capture precise vibration patterns within the target task to enhance prediction accuracy.
@@ II-A Meta-level Training on Auxiliary Tasks
To distill knowledge from auxiliary tasks and leverage vibration patterns, we enhance the MAML framework by incorporating long short-term memory (LSTM) [27] for meta-learning updates. Notably, we propose a related task-aware meta-learning approach combined with a curriculum strategy, enabling knowledge transfer by evaluating the relevance of patterns across tasks.
@@ II-A1 Basic Theory of MAML
As a subset of meta-learning, the MAML algorithm differs from traditional deep learning by focusing on training a model’s initialization parameters within its outer loop. By optimizing these initialization parameters to be broadly applicable, the model can adapt to new tasks with limited data efficiently, requiring only a few gradient updates. From a learning perspective, MAML aims to uncover general internal representations that promote effective transferability across diverse tasks. As a result, minimal parameter adjustments are needed when tackling new tasks.
Following [17], fault diagnosis across various working conditions can be framed as a single task within the meta-learning framework. Accordingly, the vibration signals from auxiliary tasks $\{T^{\mu}\}{\mu=1}^{A}$ are represented as $\mathbb{D}_{meta}^{(aux)}$, while the vibration signals from target tasks $T^{tar}$ are partitioned into training sets $\mathbb{D}_{train}^{(tar)}$ and testing sets $\mathbb{D}_{test}^{(tar)}$. Each working condition $y_{m}$ is treated as an individual meta-learning task, with a support set $\mathcal{D}^{spt}_{y_{m}}$ for training and a query set $\mathcal{D}^{qry}_{y_{m}}$ for evaluation. Ultimately, the objective is to utilize the auxiliary tasks to train a meta-learner $F_{w}$ that encapsulates knowledge across various working conditions.
Specifically, in each iteration of MAML, the process involves a local update and a global update on a batch of sampled tasks. During the local update phase, the model parameters $\theta$ are adjusted on the support set $\mathcal{D}^{spt}$ for each task. Subsequently, in the global update phase, $\theta$ is updated using gradient descent to minimize the aggregate loss computed on the query set $\mathcal{D}^{qry}$ across all tasks.
  • Local Update: A batch of tasks is first sampled from $\mathbb{D}^{(aux)}_{meta}$. For each selected task, a subset of vibration signals is randomly drawn from its support set $\mathcal{D}^{spt}_{y_{m}}$ and query set $\mathcal{D}^{qry}_{y_{m}}$. The training loss is then computed on $\mathcal{D}^{spt}_{y_{m}}$, and the parameters $\theta$ are locally updated with a single gradient descent step:
@@
  • Global Update: The testing loss is evaluated on each $\mathcal{D}^{qry}_{y_{m}}$ using the corresponding locally updated parameters $\theta_{y_{m}}^{\prime}$. The goal of the MAML global update is to update the initialization $\theta$ by performing a single gradient descent step on the aggregated testing losses across all tasks:
@@
The optimal parameters for the meta-learning model are achieved through iterative optimization within the inner (local update) and outer (global update) loops. Consequently, the objective of meta-optimization is to minimize the task-specific loss function: